### v2

# Fine-Grained Reward Optimization for Machine Translation using Error Severity Mappings

Abstract
Reinforcement learning (RL) has been proven to be an effective and robust method for training neural machine translation systems, especially when paired with powerful reward models that accurately assess translation quality. However, most research has focused on RL methods that use sentence-level feedback, leading to inefficient learning signals due to the reward sparsity problem – the model receives a single score for the entire sentence. To address this, we propose a novel approach that leverages fine-grained, token-level quality assessments along with error severity levels using RL methods. Specifically, we use xCOMET, a state-of-the-art quality estimation system, as our token-level reward model. We conduct experiments on small and large translation datasets with standard encoder-decoder and large language models-based machine translation systems, comparing the impact of sentence-level versus fine-grained reward signals on translation quality. Our results show that training with token-level rewards improves translation quality across language pairs over baselines according to both automatic and human evaluation. Furthermore, token-level reward optimization improves training stability, evidenced by a steady increase in mean rewards over training epochs.
1 Introduction
Neural machine translation (NMT) Kalchbrenner and Blunsom (2013); Sutskever et al. (2014); Cho et al. (2014), a leading approach within MT, leverages neural networks to automate language translation and has driven significant improvements in translation quality. However, most NMT systems are predominantly trained using maximum likelihood estimation (MLE). MLE-based training focuses on maximizing the probability of next-word predictions given a partial reference. This often leads to a critical problem known as exposure bias – the model uses ground-truth prefix tokens during training, but during inference it relies on its previous predictions Bengio et al. (2015); Ranzato et al. (2015); Wiseman and Rush (2016). This can cause errors to propagate through the generated sequence, severely degrading the translation quality. Furthermore, it tends to produce translations that lack global coherence and adequacy as the model does not sufficiently consider the context of entire sentences or the overarching meaning. This has spurred interest in using alternative approaches that leverage RL methods for training NMT systems.
RL-based approaches use explicit reward models to evaluate the outputs generated by the NMT system, assigning scores to generated hypotheses to guide the learning process. However, most prior research Ranzato et al. (2015); Wu et al. (2016); Bahdanau et al. (2016); Nguyen et al. (2017); Wu et al. (2017); Kreutzer et al. (2018a, b); Kiegeland and Kreutzer (2021) predominantly relies on sentence-level feedback and often struggles with reward sparsity, particularly for long-form text generation: sentence-level rewards fail to capture specific issues within a translation, making it difficult for the model to learn from negative reward signals. As shown in Figure 1, two translations corresponding to different source texts of varying length receive the same sentence-level quality score of 70, yet differ significantly in the nature and impact of the errors: The first translation has several minor errors scattered throughout the text, while the latter has major errors that could potentially hinder the understanding of the original content. This suggests that learning can be more effective if feedback is provided at a fine-grained level, including precise identification of the nature of errors.
Recent advancements in automated MT evaluation metrics that generate fine-grained error span predictions, such as xCOMET Guerreiro et al. (2024), MetricX Juraska et al. (2023), AutoMQM Fernandes et al. (2023), and MaTESE Perrella et al. (2022), have shown promise in improving alignment with human translation quality judgments. These metrics directly predict token-level error severity (no error, minor, major, or critical) and optionally sentence-level quality assessments, or prompt large language models to identify error types (e.g. mistranslation, omission) and severities based on the Multidimensional Quality Metrics (MQM) framework (Lommel et al., 2014).
Despite the potential of the severity-based metrics to improve translation quality, their application in MT training via RL methods remains relatively underexplored, since it presents several challenges: (i) the feedback, albeit informative and frequent, can be noisy, and (ii) determining the appropriate reward assignments for different severity levels to ensure effective and stable learning is not straightforward. In this regard, our research aims to answer the following questions:
1.
Does fine-grained RL methods offer benefit over sentence-level feedback in improving translation quality and stabilizing training?
2.
Can automatic fine-grained MT metrics be effectively used to provide accurate, detailed, human-aligned feedback to reduce reward sparsity?
When answering these questions, we make the following contributions:
1.
We propose using a fine-grained evaluation metric, xCOMET, to generate token-level rewards, which increases the reward density by providing frequent token-level rewards, thus improving the robustness and stability of RL-based MT.
2.
We introduce a new severity map to effectively use the reward signals, overcoming the limitations of standard MQM scoring, as demonstrated in our experimental results.
3.
We conduct experiments on English-to-German (ENDE), English-to-French (ENFR), German-to-English (DEEN), and French-to-English (FREN) translation datasets, comparing the overall translation quality of NMT systems when using sentence and token-level rewards, showing that translation quality improves when employing xCOMET as a reward model.
By integrating fine-grained reward signals into NMT training, we demonstrate significant improvements in translation quality and overcome the challenges of exposure bias, reward sparsity, and stability of RL training, paving the way for more reliable and accurate MT systems.
2 Background
Standard NMT Training.
NMT systems utilize learnable parameters, denoted as , to estimate the probability distribution over a set of possible translations , conditioned on a given source sentence . In the simplest form of NMT training, maximum likelihood estimation (MLE) is used, which maximizes the probability of the correct target translation given the source sentence . The MLE objective can be expressed as:
where represents a dataset of parallel sentences.
Limitations of MLE Training.
While commonly used in NMT, MLE has several limitations, primarily its weak learning signals from token-level feedback. As MLE assumes gold-reference tokens (teacher-forcing) during training, when exposed to its own incorrect predictions during inference, it can lead to error accumulation and poor performance on longer sequences. Another major limitation is its tendency to optimize for a single “most likely” translation, often ignoring the variety of equally valid alternatives, which reduces the model’s ability to generate diverse and natural outputs. Additionally, MLE is sensitive to noisy or inconsistent reference translations, which can degrade performance by producing unreliable gradient updates. Taken together, these challenges have prompted the exploration of RL methods, which offer more effective feedback on model-generated outputs by optimizing directly for downstream translation quality measures.
Formulating MT as an RL Problem.
In the context of MT, we can model the translation process as a Markov Decision Process (MDP) Puterman (1990), defined by the tuple with a finite vocabulary . The state space consists of all possible sequences of tokens up to the current time step, which includes the input sequence in the source language as well as the target language tokens generated so far. Initially, the state corresponds to the input sentence in the source language, , where each token . At each time step , the state represents the sequence of tokens generated up to that point, which can be expressed as:
The agent selects an action , which is a token generated by the policy based on the current state . The process continues until an end-of-sequence token is generated, completing the translation. The reference tokens in the target language are denoted by , where . The generated tokens are evaluated against to measure the quality of the translation. For , the state transition function defines the probability of transitioning from one state to another by appending a chosen token to the current translation, and the reward function assigns a real-valued reward to each transition , where and , based on the quality of the generated translation sequence. Explicitly, the reward function can be written as:
Sequence-level rewards are provided only once per translation and evaluate the entire output at once. Token-level rewards, on the other hand, give feedback for each generated token. The discount factor is used to weigh future rewards, with typically chosen in MT to ensure that all rewards are valued equally, allowing the optimization of the entire sequence of tokens in the translation rather than focusing on just the initial tokens. Finally, the goal is to maximize the expected cumulative reward over trajectories sampled from the . The objective function can be written as:
Policy Gradient Algorithms
To optimize the above objective, we can use policy gradient methods: REINFORCE Williams (1992); Ranzato et al. (2015), a vanilla policy gradient method, optimizes translation by sampling hypotheses , scoring them with a reward obtained from an MT metric , and updating the model to maximize expected rewards, as shown in Equation 2. Despite its simplicity, it often struggles with high variance and instability. Proximal Policy Optimization (PPO) Schulman et al. (2017) mitigates this by using a clipped surrogate objective (Equation 3) to keep policy updates stable within a margin , efficient and employs Generalized Advantage Estimation (GAE) Schulman et al. (2015) to compute advantages using rewards and value function . While PPO performs well across various tasks, simpler methods like REINFORCE can sometimes rival or surpass it Ahmadian et al. (2024). Both are evaluated in our experiments (§5.2).
3 Related Work
Advancements and Challenges in Sentence-level Feedback.
Incorporating human feedback as rewards and optimizing language models with RL methods effectively aligns them with human preferences Ouyang et al. (2022), often surpassing MLE. A notable example in translation tasks is Minimum Risk Training (MRT) Shen et al. (2016), which minimizes expected risk based on evaluation metrics to directly improve translation quality. Recent advances in NMT build on this idea by refining training with feedback from metrics or human evaluations, incorporating alignment techniques and RL methods Nguyen et al. (2017); Kreutzer et al. (2018b); Wu et al. (2018); Kiegeland and Kreutzer (2021); Xu et al. (2023); Agrawal et al. (2024); Zhu et al. (2024); He et al. (2024); Ramos et al. (2024). Despite these advancements, sentence-level feedback methods face persistent challenges such as sparse rewards, instability, and difficulty handling long sequences Wu et al. (2018). These issues hinder performance, generalization, and robust learning, even with multi-objective optimizations Wu et al. (2023); Jang et al. (2023). To address the limitations of sentence-level feedback, recent research has explored finer-grained rewards at the token level for tasks such as language model alignment Xia et al. (2024); Yoon et al. (2024); Guo et al. (2023); Cao et al. (2024), controllable text generation Li et al. (2024), query generation Ouyang et al. (2024), among others, but remain relatively underexplored in MT.
Token-level Feedback and Reward Modeling for MT.
Previous approaches to token-level reward modeling often relied on binary error markings generated by humans Kreutzer et al. (2020); Domingo et al. (2017) or simulated it by comparing model predictions with reference translations based on heuristic methods Petrushkov et al. (2018). While effective, these methods provide limited feedback due to their binary nature and require costly human annotation, making them less practical for scalable solutions. Other approaches have employed reward shaping techniques Ng et al. (1999); Wu et al. (2018); Goyal et al. (2019); Devidze et al. (2022), incorporating intermediate rewards along with BLEU Papineni et al. (2002) as the reward function. However, partial BLEU or token-level BLEU are less effective for fine-grained reward modeling, as they depend on exact -gram matching and fail to capture meaningful semantic differences and context. Consequently, these methods, while valuable, are limited in their granularity and fail to address the severity of errors introduced at the token level.
To address these limitations, in this work, we instead use token-level rewards obtained from state-of-the-art evaluation metrics that predict error spans and error severity levels to improve the learning and translation quality of MT systems.
4 Approach
In this section, we present our method for incorporating token-level rewards into RL training for machine translation (MT). We first detail how error assessments are mapped to token-level reward scores and then describe our adaptation of the REINFORCE and PPO learning objectives at the token level to enable learning from this fine-grained feedback.
Token-level Reward Modeling.
Building on the MDP formulation for MT, we focus on token-level reward modeling – feedback is provided for individual tokens rather than entire sequences – allowing the model to refine its policy by identifying and addressing specific translation errors. Given an evaluation metric, that predicts error spans along with their severity levels (e.g. minor, major, critical) for a hypothesis given source and optionally a gold reference, we assign numerical weights to each token within an error span according to a severity mapping as defined below:
We use the evaluation metric, xCOMET as as it was shown to achieve the best correlation with human judgments and was the winning submission for the WMT23 Metrics Shared Task Freitag et al. (2023). The severity weights from xCOMET adhere to the MQM framework Lommel et al. (2014), which classifies translation issues into categories such as fluency, adequacy, grammar, and style. Each token within an error span is assigned the same severity weight (see Figure 1). We note that although the weights follow the MQM guidelines, they need to be further adjusted depending on the tasks to optimize the performance of token-level RL.
Tokenization-Agnostic Reward Assignment.
MT as well as evaluation models often rely on subword-level tokenization methods, such as Byte-Pair Encoding (Gage, 1994, BPE) or SentencePiece Kudo and Richardson (2018), where a single word can be split into multiple tokens. For a detokenized hypothesis generated by the MT system, identifies detokenized error spans corresponding to the errors. To generate token-level reward assignments for the tokenized hypothesis, , we assign the same reward to all subword tokens that belong to the same parent word. This is done by identifying the boundaries of the original word and ensuring that all tokens within those boundaries share the same reward. In this way, token-level rewards remain robust and independent of the tokenization differences between the generation and the evaluation models.
Token-level Policy Refinement.
In token-level RL, we maintain the structure of traditional sentence-level RL losses but adjust them to operate at the token level. We generate the full sequence, compute rewards for each token, and then perform updates for each token separately. The traditional sentence-level REINFORCE objective (Equation 2) is adapted to token-level by adjusting the loss to calculate the reward for each individual token. After generating the full sequence, we perform updates for each token one at a time, as follows:
Here, is the reward for token , reflecting its contribution to the overall sequence. Similarly, we extend the sentence-level PPO objective (Equation 3) to token-level by modifying the loss function to compute the policy ratio and advantage for each token independently. The token-level PPO objective is defined as:
The policy ratio captures the change in policy for each token relative to the previous policy. The token-level advantage measures the improvement in performance over a baseline for each token and is used to scale updates based on expected reward improvements, as described in Schulman et al. (2015, 2017). To ensure stable training, we apply clipping to limit the extent of policy updates, preventing large, unstable shifts. This approach allows for more granular control over the model’s learning, ensuring that each token is generated in a way that maximizes task-specific objectives while maintaining stability in the policy updates. Clipping also helps mitigate length bias by preventing longer sequences from accumulating disproportionately high rewards.
5 Experiments
We outline the experiments designed to explore the application of RL for MT, specifically focusing on comparing the impact of sentence-level and token-level reward signals.
5.1 Experimental Setup
Models.
We use three state-of-the-art models: a standard encoder-decoder MT model, NLLB111facebook/nllb-200-1.3B NLLB Team et al. (2022), and two LLM-based MT systems, Tower222Unbabel/TowerInstruct-Mistral-7B Alves et al. (2024) and Gemma333https://huggingface.co/google/gemma-2-9b-it Gemma Team et al. (2024). While NLLB and Tower are dedicated MT models optimized for translation tasks, Gemma is an LLM that exhibits strong multilingual capabilities. These models differ in both their architectures and pre-training methodologies. Each is pre-trained on diverse multilingual datasets, establishing them as robust baselines for investigating the effects of SFT and RL techniques.
Data.
We use the following training datasets in our experiments: (1) The IWSLT2017 dataset Cettolo et al. (2017), with k examples for English-French (ENFR), supports rapid experimentation and frequent training iterations. (2) The WMT18 dataset Bojar et al. (2018) contains M examples for English-German (ENDE). We train NLLB with both datasets and the LLM-based models with (2). Training stops once rewards stabilize, so not all examples are used.
We evaluate NLLB models using the IWSLT17 ENFR and WMT18 ENDE test sets. To standardize comparison across MT systems (NLLB and LLM-based MT systems), we also evaluate all models on the WMT24 dataset Kocmi et al. (2024), addressing concerns about data contamination, as Tower training included the WMT18 test set.
Evaluation.
We assess translation quality using a comprehensive suite of well-established evaluation metrics. These include lexical reference-based metrics, such as BLEU Papineni et al. (2002) and ChrF Popović (2015); neural reference-based metrics, including COMET22 Rei et al. (2022), xCOMET Guerreiro et al. (2024), and BLEURT Sellam et al. (2020); and a neural reference-free metric, CometKiwi-23 Rei et al. (2023). Lexical metrics focus on word overlap and -gram matching, while neural metrics evaluate translations in terms of semantic coherence and contextual quality. Including a reference-free metric enables evaluation without reliance on predefined reference texts. This diverse set of metrics captures multiple dimensions of translation quality, including fluency, grammatical accuracy, semantic adequacy, and contextual relevance. By using various evaluation criteria, we reduce potential biases that may arise from aligning the reward model with a single evaluation metric, ensuring more robust and reliable conclusions about the impact of different approaches on translation quality.
We apply significance testing at a confidence threshold of . For segment-level metrics, such as COMET-22, we test at the segment level, but for corpus-level metrics, such as BLEU and ChrF, we apply bootstrapping with samples of size Koehn (2004). Performance clusters are formed based on statistically significant gaps, and final rankings are derived by averaging the cluster scores across all languages Colombo et al. (2022); Freitag et al. (2023).
In addition to automated metrics, we conduct human evaluations to gain deeper insights into translation performance. Furthermore, significance testing is performed to validate improvements observed across different models and approaches, ensuring the reliability of our findings.
Reward Models.
We utilize two reward models based on xCOMET444https://huggingface.co/Unbabel/XCOMET-XL. The MQM-derived reward signal, referred to as xCOMET-MQM, is generated from error span predictions that identify and classify translation errors by severity. Token-level reward signals are directly computed from these error spans, while sentence-level rewards are obtained as a weighted average of the token-level severity spans. We also use the standard sentence-level reward signal provided by xCOMET as a baseline for comparison.
Training Configurations.
We finetune NLLB, Tower and Gemma models using MLE and RL methods detailed in §3 with the following configurations:
SFT: a baseline model supervised finetuned on the parallel data using MLE (1).
sRL: We compare using sentence-level xCOMET with BLEU. The learning algorithm used is PPO (3), a current state-of-the-art alignment method for MT.
CPO Xu et al. (2024): a state-of-the-art preference optimization learning method for MT, offering a more efficient variant of DPO Rafailov et al. (2023). We construct the preference dataset by generating multiple outputs from the MT model using the training datasets,555We generate 16 samples with the value of top_p top_k set to 0.9 and 50 respectively. and then induce preferences using the xCOMET metric, comparing these outputs to human-written references.
Hyperparameter Details.
We use the same hyperparameter settings for NLLB, Tower, and Gemma. We use HuggingFace’s Transformers library Wolf et al. (2020) and the Transformers Reinforcement Learning (TRL) library 666https://github.com/huggingface/trl to facilitate RL training. We perform MLE training with Adam Kingma and Ba (2014) as the optimization algorithm, learning rate decay starting from and early stopping. We use PPO with a learning rate of , set as , trajectory limit set as . Mini-batch updates are performed with a batch size of over PPO epochs. The translation prompt for LLM-based MT systems is shown below:
| Translate the following text from {source_lang} into {target_lang}. |
| {source_lang}: {source_sentence}. |
| {target_lang}: |
5.2 Results and Main Findings
| Model | Metrics | |||||
| BLEU | ChrF | COMET22 | xCOMET | BLEURT | CometKiwi-23 | |
| WMT18 EN→DE | ||||||
| NLLB | 41.60 | 66.34 | 86.94 | 94.69 | 76.42 | 72.46 |
| + SFT | 43.07 | 67.38 | 87.18 | 94.97 | 76.35 | 72.57 |
| + sRL w/ BLEU | 43.31 | 67.57 | 87.31 | 95.13 | 76.61 | 72.73 |
| + sRL w/ xCOMET | 43.09 | 67.34 | 87.18 | 95.55 | 76.72 | 72.84 |
| + sRL w/ xCOMET-MQM | 43.03 | 67.29 | 87.10 | 95.50 | 76.72 | 72.40 |
| + tRL w/ BLEU | 42.90 | 67.32 | 87.15 | 94.94 | 76.37 | 72.57 |
| + tRL w/ xCOMET-MQM | 42.63 | 67.47 | 88.27 | 96.18 | 77.60 | 75.16 |
| + CPO w/ xCOMET | 42.30 | 66.10 | 87.02 | 95.31 | 76.70 | 73.08 |
| WMT18 DE→EN | ||||||
| NLLB | 43.50 | 66.12 | 86.45 | 93.95 | 75.84 | 73.16 |
| + SFT | 45.57 | 67.83 | 87.22 | 95.08 | 76.79 | 74.11 |
| + sRL w/ BLEU | 45.78 | 67.97 | 87.26 | 95.07 | 76.85 | 74.05 |
| + sRL w/ xCOMET | 45.51 | 67.78 | 87.21 | 95.28 | 76.76 | 74.06 |
| + sRL w/ xCOMET-MQM | 45.55 | 67.80 | 87.21 | 95.17 | 76.80 | 74.11 |
| + tRL w/ BLEU | 45.53 | 67.74 | 87.29 | 95.14 | 76.84 | 74.00 |
| + tRL w/ xCOMET-MQM | 45.71 | 67.60 | 87.91 | 96.02 | 77.70 | 74.42 |
| + CPO w/ xCOMET | 43.88 | 66.05 | 87.10 | 95.33 | 77.08 | 74.23 |
| IWSLT2017 EN→FR | ||||||
| NLLB | 43.61 | 65.76 | 84.93 | 90.27 | 72.61 | 70.06 |
| + SFT | 45.63 | 67.59 | 86.01 | 91.07 | 74.41 | 71.78 |
| + sRL w/ BLEU | 45.77 | 67.68 | 86.04 | 91.11 | 74.44 | 71.79 |
| + sRL w/ xCOMET | 45.68 | 67.61 | 86.03 | 91.67 | 74.45 | 71.77 |
| + sRL w/ xCOMET-MQM | 45.76 | 67.60 | 86.08 | 91.68 | 74.40 | 71.77 |
| + tRL w/ BLEU | 45.71 | 67.63 | 86.00 | 91.14 | 74.48 | 71.82 |
| + tRL w/ xCOMET-MQM | 46.58 | 60.07 | 87.17 | 92.17 | 75.58 | 72.80 |
| + CPO w/ xCOMET | 43.96 | 65.24 | 85.55 | 91.65 | 74.02 | 71.86 |
| IWSLT2017 FR→EN | ||||||
| NLLB | 45.76 | 65.76 | 87.15 | 94.69 | 77.55 | 71.99 |
| + SFT | 48.65 | 67.69 | 88.22 | 95.34 | 78.74 | 73.59 |
| + sRL w/ BLEU | 48.76 | 67.76 | 88.22 | 95.33 | 78.74 | 73.54 |
| + sRL w/ xCOMET | 48.62 | 67.88 | 88.30 | 95.53 | 78.73 | 73.56 |
| + sRL w/ xCOMET-MQM | 48.61 | 67.66 | 88.31 | 95.53 | 78.74 | 73.56 |
| + tRL w/ BLEU | 48.56 | 67.69 | 88.21 | 95.33 | 78.73 | 73.60 |
| + tRL w/ xCOMET-MQM | 49.06 | 61.17 | 88.46 | 95.27 | 79.10 | 74.74 |
| + CPO w/ xCOMET | 48.46 | 67.54 | 88.20 | 95.42 | 78.71 | 73.52 |
We present the main results of comparing the different methods across datasets trained using NLLB in Table 2 and across models in Table 3.
tRL consistently outperforms SFT and sRL methods across neural metrics.
For all language directions reported in Table 2, “tRL w/ xCOMET-MQM“ outperforms SFT and its sentence-level counterpart “sRL w/ xCOMET-MQM“ across all neural metrics considered. SFT significantly improves translation quality by tailoring the pre-trained MT model to the specific target language pairs. Moreover, applying RL methods (sRL or tRL) on top of SFT further enhances the MT model’s performance by directly optimizing translations based on targeted reward signals. When comparing the sRL and tRL methods, we observe that sRL leads to moderate improvements over SFT, while the gains obtained by tRL are more substantial, particularly when assessed with advanced neural metrics. Although the ChrF scores for tRL models trained with xCOMET-MQM are lower, possibly due to the focus on optimizing token-level features over character-level accuracy, all other metrics indicate significant improvements, suggesting the overall translation quality is better.
| Model | Metrics | |||||
| BLEU | ChrF | COMET22 | xCOMET | BLEURT | CometKiwi-23 | |
| WMT24 EN→DE | ||||||
| NLLB | 35.01 | 53.92 | 64.30 | 83.40 | 60.90 | 55.50 |
| + SFT | 37.92 | 59.29 | 65.20 | 82.50 | 61.41 | 58.20 |
| + sRL w/ BLEU | 38.54 | 59.77 | 65.50 | 82.50 | 61.40 | 58.30 |
| + sRL w/ xCOMET | 37.88 | 59.16 | 65.50 | 82.80 | 60.40 | 57.90 |
| + sRL w/ xCOMET-MQM | 37.14 | 59.29 | 65.40 | 83.30 | 60.40 | 57.50 |
| + tRL w/ BLEU | 38.53 | 59.60 | 65.49 | 82.80 | 61.41 | 59.10 |
| + tRL w/ xCOMET-MQM | 39.42 | 60.40 | 66.43 | 83.60 | 62.29 | 60.70 |
| + CPO w/ xCOMET | 36.07 | 55.32 | 64.70 | 82.90 | 60.80 | 58.10 |
| Tower | 42.77 | 62.60 | 71.80 | 88.50 | 68.59 | 67.20 |
| + SFT | 45.14 | 63.30 | 72.18 | 88.90 | 69.18 | 67.30 |
| + sRL w/ BLEU | 46.07 | 63.77 | 72.26 | 89.20 | 70.01 | 68.20 |
| + sRL w/ xCOMET | 45.59 | 64.11 | 72.73 | 90.50 | 70.58 | 69.90 |
| + sRL w/ xCOMET-MQM | 45.58 | 63.76 | 72.41 | 90.20 | 70.55 | 69.70 |
| + tRL w/ BLEU | 46.13 | 64.15 | 72.22 | 89.22 | 70.11 | 68.33 |
| + tRL w/ xCOMET-MQM | 46.92 | 65.63 | 74.66 | 91.90 | 71.80 | 71.20 |
| + CPO w/ xCOMET | 43.15 | 61.13 | 70.65 | 87.80 | 68.23 | 67.38 |
| Gemma | 15.13 | 51.99 | 54.84 | 61.40 | 51.35 | 42.80 |
| + SFT | 35.19 | 59.03 | 69.86 | 86.10 | 66.13 | 64.30 |
| + sRL w/ BLEU | 36.34 | 59.33 | 69.96 | 86.10 | 66.50 | 64.40 |
| + sRL w/ xCOMET | 36.16 | 59.28 | 70.82 | 87.20 | 66.90 | 66.30 |
| + sRL w/ xCOMET-MQM | 35.98 | 59.33 | 70.39 | 87.10 | 67.20 | 65.70 |
| + tRL w/ BLEU | 36.63 | 59.52 | 70.02 | 86.40 | 66.70 | 64.50 |
| + tRL w/ xCOMET-MQM | 36.90 | 59.58 | 71.12 | 88.10 | 68.00 | 66.90 |
| + CPO w/ xCOMET | 35.05 | 58.52 | 69.22 | 86.40 | 66.80 | 65.10 |
tRL improves translation quality for LLM-based MT systems, Tower and Gemma.
Our severity-based fine-grained reward optimization mechanism significantly improves translation quality across all automatic evaluation metrics as shown in Table 3. These findings highlight that tRL not only improves the quality of state-of-the-art MT models but can also significantly boost stronger MT-LLMs, demonstrating its broad applicability and potential for advancing multilingual translation systems.
On-policy PPO results in better translation quality than RL-free method, CPO.
Both sentence-level and token-level RL methods achieve higher evaluation scores than CPO, demonstrating significant improvement in translation quality across language pairs. Unlike CPO, which focuses on maintaining predefined constraints imposed by the preference dataset, RL methods like PPO can flexibly and dynamically adjust the MT model based on real-time feedback from the reward models via iterative feedback and refinement. Furthermore, tRL uses fine-grained reward signals from xCOMET-MQM capturing a wider range of linguistic features and quality indicators, thus offering more precise and contextually relevant feedback during the training process. This feedback can be leveraged more effectively with PPO than with CPO.
xCOMET is a superior reward model than lexical MT metrics.
The role of the reward model in achieving alignment is crucial, as also evidenced by our findings (Tables 2 and 3). Our results clearly show that using xCOMET as a reward model, particularly at the token level, significantly improves translation quality as measured by several metrics. Given that xCOMET exhibits a strong correlation with human judgments, it proves to be an essential tool for guiding MT models toward higher translation quality. In contrast, traditional metrics like BLEU, based on -gram overlap, can fall short in aligning with human judgments as they do not capture contextual nuances and semantic understanding Freitag et al. (2022). Consequently, BLEU performs less effectively compared to neural metrics like xCOMET in this setup which use contextual embeddings. Therefore, incorporating neural metrics as reward models is crucial for capturing the subtleties of language and improving the overall quality and reliability of MT models.
5.3 Ablation Study
We present several ablations to study how the design choices employed impact the learning and the final translation quality of the optimized model.
| Word | Bin | MQM | rMQM | Our | rOur |
| Correct | 1 | 0 | 25 | 8 | -1 |
| Minor | -1 | -1 | 5 | 4 | -2 |
| Major | -1 | -5 | 1 | 2 | -4 |
| Critical | -1 | -25 | 0 | 1 | -8 |
| Model | Severity Map | Metrics | |||||
| BLEU | ChrF | COMET22 | xCOMET | BLEURT | CometKiwi-23 | ||
| IWSLT2017 EN→FR | |||||||
| NLLB | – | 43.61 | 65.76 | 84.93 | 90.27 | 72.61 | 70.06 |
| + SFT | – | 45.63 | 67.59 | 86.01 | 91.07 | 74.41 | 71.78 |
| + tRL w/ xCOMET-MQM | Bin | 45.44 | 60.43 | 85.87 | 90.90 | 74.23 | 71.58 |
| + tRL w/ xCOMET-MQM | MQM | 45.68 | 60.67 | 85.97 | 91.01 | 74.33 | 71.74 |
| + tRL w/ xCOMET-MQM | rMQM | 41.47 | 59.74 | 82.20 | 84.58 | 70.45 | 67.29 |
| + tRL w/ xCOMET-MQM | Our | 46.58 | 60.07 | 87.17 | 92.17 | 75.58 | 72.80 |
| + tRL w/ xCOMET-MQM | rOur | 45.89 | 60.10 | 85.90 | 91.33 | 74.39 | 71.79 |
Choice of severity map impacts learning.
We investigate the impact of different severity maps on token-level RL training using xCOMET, as detailed in Table 4. The severity maps we evaluate include the default MQM-based map (MQM), our custom map (Our), the reversed MQM-based map (rMQM), the reversed custom map (rOur), and a binary map (Bin). Our findings, shown in Table 5, highlight the importance of having gradual transitions between reward values. Smooth transition severity map to result in better translation quality. In contrast, abrupt changes in the reward signal can destabilize learning, leading to inconsistent training, oscillations, or convergence to suboptimal policies Ranzato et al. (2015); Sutton and Barto (2018). Additionally, we find that the binary severity map, which ignores the severity of errors, provides less informative feedback to the model, resulting in slightly lower performance than maps that offer more nuanced assessments. Although designing custom severity maps can increase complexity and require hyperparameter tuning, our experiments suggest they can be set in a straightforward way with minimal overhead. We leave a more systematic investigation of these mappings, including the possibility of learning them during training, to future work.
tRL improves training stability over sRL.
Figure 3 shows the evolution of mean rewards during training for sRL and tRL methods across the two datasets for the NLLB system. As can be observed, tRL training shows a more stable and consistently increasing reward evolution, which is crucial for the training process as it ensures consistent improvements and reduces the risk of overfitting or fluctuations that could degrade the MT model’s performance.
tRL improves translation quality for longer sequences.
Building on our hypothesis that tRL is particularly effective for longer sequences, we show the COMET22 scores for the WMT24 EN DE dataset, grouped by sequence length, across different training methods, including NLLB and Tower, in Figure 4. tRL consistently outperforms other methods, especially for longer sequences, which we attribute to its ability to capture localized and fine-grained signals. These results further support our findings, showing that tRL remains the most robust method, with scores dropping less than other approaches as the source sentence length increases. This reinforces the effectiveness of tRL in handling complex, longer sentence structures.
| Model | Metrics | |||||
| BLEU | ChrF | COMET22 | xCOMET | BLEURT | CometKiwi-23 | |
| IWSLT2017 EN→FR | ||||||
| NLLB | 43.61 | 65.76 | 84.93 | 90.27 | 72.61 | 70.06 |
| + SFT | 45.63 | 67.59 | 86.01 | 91.07 | 74.41 | 71.78 |
| + sREINFORCE w/ xCOMET | 45.70 | 67.58 | 86.11 | 91.29 | 74.43 | 71.75 |
| + sREINFORCE w/ xCOMET-MQM | 45.72 | 67.59 | 86.07 | 91.35 | 74.78 | 71.80 |
| + tREINFORCE w/ xCOMET-MQM | 46.07 | 60.60 | 86.24 | 91.77 | 74.98 | 71.87 |
| + sPPO w/ xCOMET | 45.68 | 67.61 | 86.03 | 91.67 | 74.45 | 71.77 |
| + sPPO w/ xCOMET-MQM | 45.76 | 67.60 | 86.08 | 91.68 | 74.40 | 71.77 |
| + tPPO w/ xCOMET-MQM | 46.58 | 60.07 | 87.17 | 92.17 | 75.58 | 72.80 |
REINFORCE and PPO are suitable methods for training MT systems.
We compare REINFORCE and PPO for RL-based MT with xCOMET as reward model, evaluating their impact on translation quality (Table 6). Both methods are effective, but PPO achieves superior overall metric scores due to features such as objective clipping and KL divergence control, which enhance training stability. However, REINFORCE remains a strong alternative for simpler implementations that aim to achieve competitive performance.
| Reward Model | BLEU | xCOMET |
| Latency | ||
| Throughput | ||
| Quality |
Efficiency and Quality Tradeoffs in Token-Level Reward Computation.
tRL using xCOMET offers higher quality, as measured by COMET22, than tRL with BLEU but comes with increased computational costs, including higher latency and lower throughput due to its larger pre-trained encoder, as reported in Table 7. Despite the slower processing speeds, xCOMET’s superior reward quality and resulting higher model quality are particularly valuable in token-level feedback scenarios, where high-quality rewards are essential, and the additional computational cost is manageable with GPU acceleration. Furthermore, advances in compute efficiency for transformer-based architectures, such as quantization Jacob et al. (2017), FlashAttention Dao et al. (2022); Dao (2023), and distillation Hinton et al. (2015), can help reduce the computational load, making xCOMET more practical for broader applications.
5.4 Human Evaluation
Setup.
For our human evaluation, we used Direct Assessments (Graham et al., 2013, DAs) to score translations on a scale from 0 to 100, following the standard WMT human evaluation methodology. We evaluate 200 randomly chosen instances from the WMT18 EN DE dataset. Two professional translators, both native speakers of the target language, assesse the references and NLLB translations using the following methods: SFT, CPO with xCOMET, tRL with xCOMET and our proposed severity mapping, and sRL with xCOMET.
While WMT18 EN DE allows for easy comparison between several methods due to shorter sequences, we conducte a second human evaluation on the WMT24 EN DE dataset to validate our empirical finding that tRL benefits longer input sequences. For this setting, we directly compare sRL with xCOMET and tRL with xCOMET in a pairwise setting using outputs from the Tower model. To ensure adequate coverage across different sequence lengths, we performed stratified sampling based on source length, using bins [0, 100, 250, 500, 1000] with 50 instances per bin.
| Model | Ann1 | Ann2 |
| CPO | 59.5 | 53.2 |
| SFT | 64.5 | 54.2 |
| tRL | 66.9 | 56.1 |
| sRL | 66.8 | 57.6 |
| Reference | 76.0 | 60.0 |
Findings.
As shown in Table 8, both sRL and tRL models consistently outperform SFT and CPO, demonstrating their advantage in enhancing translation quality. On average, human evaluation scores for tRL and sRL are similar. On the WMT24 dataset, tRL achieves an average DA score of 82.6, 2.3 points higher than sRL. We additionally show the distribution of DA scores bucketed by length in Figure 5. Across all buckets, tRL outperforms sRL, with more significant (based on error bars overlap) gains for longer sentences. These findings are also consistent with our automatic evaluation, including the ablation analysis (§ 5.3) and clearly demonstrate that tRL enhances stability and translation quality, particularly for longer sentences.
6 Conclusion
In this work, we propose a new method for improving NMT that uses fine-grained reward optimization with xCOMET as a token-level reward model. While exposure bias arises from SFT, sentence-level RL addresses this issue but introduces reward sparsity due to coarse-grained feedback. Our token-level RL approach overcomes this by providing a denser and more informative reward signal to enhance translation quality. Our experiments show that incorporating fine-grained reward mechanisms significantly improves MT quality, especially for longer sequences, and also stabilizes training. Additionally, token-level RL training outperforms sentence-level RL training in most evaluation metrics. Our findings show that fine-grained RL offers a more effective MT optimization framework by mitigating reward sparsity and aligning better with human judgments.
Acknowledgments
We thank the members of SARDINE lab for their useful and constructive comments. This work was supported by the Portuguese Recovery and Resilience Plan through project C645008882-00000055 (Center for Responsible AI), by the EU’s Horizon Europe Research and Innovation Actions (UTTER, contract 101070631), by the project DECOLLAGE (ERC-2022-CoG 101088763), and by Fundação para a Ciência e Tecnologia through contract UIDB/50008/2020.

### v3

@@
 1 Introduction
-Neural machine translation (NMT) Kalchbrenner and Blunsom (2013); Sutskever et al.
-(2014); Cho et al.
-(2014), a leading approach within MT, leverages neural networks to automate language translation and has driven significant improvements in translation quality.
+Neural machine translation (NMT) Kalchbrenner2013RecurrentCT; sutskever2014sequence; cho2014properties, a leading approach within MT, leverages neural networks to automate language translation and has driven significant improvements in translation quality.
 However, most NMT systems are predominantly trained using maximum likelihood estimation (MLE).
 MLE-based training focuses on maximizing the probability of next-word predictions given a partial reference.
-This often leads to a critical problem known as exposure bias – the model uses ground-truth prefix tokens during training, but during inference it relies on its previous predictions Bengio et al.
-(2015); Ranzato et al.
-(2015); Wiseman and Rush (2016).
+This often leads to a critical problem known as exposure bias – the model uses ground-truth prefix tokens during training, but during inference it relies on its previous predictions BENGIO2015; RANZATO2016; WISEMAN2016.
 This can cause errors to propagate through the generated sequence, severely degrading the translation quality.
@@
 RL-based approaches use explicit reward models to evaluate the outputs generated by the NMT system, assigning scores to generated hypotheses to guide the learning process.
-However, most prior research Ranzato et al.
-(2015); Wu et al.
-(2016); Bahdanau et al.
-(2016); Nguyen et al.
-(2017); Wu et al.
-(2017); Kreutzer et al.
-(2018a, b); Kiegeland and Kreutzer (2021) predominantly relies on sentence-level feedback and often struggles with reward sparsity, particularly for long-form text generation: sentence-level rewards fail to capture specific issues within a translation, making it difficult for the model to learn from negative reward signals.
+However, most prior research RANZATO2016; wu2016google; Bahdanau2016; nguyen-etal-2017-reinforcement; wuseq2017; kreutzer-etal-2018-neural; kreutzer-etal-2018-reliability; kiegeland-kreutzer-2021-revisiting predominantly relies on sentence-level feedback and often struggles with reward sparsity, particularly for long-form text generation: sentence-level rewards fail to capture specific issues within a translation, making it difficult for the model to learn from negative reward signals.
 As shown in Figure 1, two translations corresponding to different source texts of varying length receive the same sentence-level quality score of 70, yet differ significantly in the nature and impact of the errors: The first translation has several minor errors scattered throughout the text, while the latter has major errors that could potentially hinder the understanding of the original content.
 This suggests that learning can be more effective if feedback is provided at a fine-grained level, including precise identification of the nature of errors.
-Recent advancements in automated MT evaluation metrics that generate fine-grained error span predictions, such as xCOMET Guerreiro et al.
-(2024), MetricX Juraska et al.
-(2023), AutoMQM Fernandes et al.
-(2023), and MaTESE Perrella et al.
-(2022), have shown promise in improving alignment with human translation quality judgments.
-These metrics directly predict token-level error severity (no error, minor, major, or critical) and optionally sentence-level quality assessments, or prompt large language models to identify error types (e.g.
-mistranslation, omission) and severities based on the Multidimensional Quality Metrics (MQM) framework (Lommel et al., 2014).
-Despite the potential of the severity-based metrics to improve translation quality, their application in MT training via RL methods remains relatively underexplored, since it presents several challenges: (i) the feedback, albeit informative and frequent, can be noisy, and (ii) determining the appropriate reward assignments for different severity levels to ensure effective and stable learning is not straightforward.
+Recent advancements in automated MT evaluation metrics that generate fine-grained error span predictions, such as xCOMET xcomet, MetricX juraska-etal-2023-metricx, AutoMQM fernandes-etal-2023-devil, EAPrompt (lu-etal-2024-error), MaTESE perrella-etal-2022-matese, and BARTScore++ (lu-etal-2023-toward) have shown promise in improving alignment with human translation quality judgments.
+These metrics directly predict token-level error severity (no error, minor, major, or critical) and optionally provide sentence-level quality assessments or prompt large language models to identify error types (e.g., mistranslation, omission) and severities based on the Multidimensional Quality Metrics (MQM) framework (lommel-mqm).
+Despite the potential of severity-based metrics to improve translation quality, their application in MT training via RL methods remains relatively underexplored, since it presents several challenges: (i) the feedback, albeit informative and frequent, can be noisy, and (ii) determining the appropriate reward assignments for different severity levels to ensure effective and stable learning is not straightforward.
 In this regard, our research aims to answer the following questions:
@@
 2.
-Can automatic fine-grained MT metrics be effectively used to provide accurate, detailed, human-aligned feedback to reduce reward sparsity?
+Can fine-grained MT metrics be effectively used to provide accurate, detailed, human-aligned feedback to reduce reward sparsity?
 When answering these questions, we make the following contributions:
@@
 We conduct experiments on English-to-German (ENDE), English-to-French (ENFR), German-to-English (DEEN), and French-to-English (FREN) translation datasets, comparing the overall translation quality of NMT systems when using sentence and token-level rewards, showing that translation quality improves when employing xCOMET as a reward model.
-By integrating fine-grained reward signals into NMT training, we demonstrate significant improvements in translation quality and overcome the challenges of exposure bias, reward sparsity, and stability of RL training, paving the way for more reliable and accurate MT systems.
+By integrating fine-grained reward signals into NMT training, we demonstrate significant improvements in translation quality and overcome the challenges of exposure bias, reward sparsity, and instability of RL training, paving the way for more reliable and accurate MT systems.
 2 Background
@@
 Formulating MT as an RL Problem.
-In the context of MT, we can model the translation process as a Markov Decision Process (MDP) Puterman (1990), defined by the tuple with a finite vocabulary .
-The state space consists of all possible sequences of tokens up to the current time step, which includes the input sequence in the source language as well as the target language tokens generated so far.
+In the context of MT, we can model the translation process as a Markov Decision Process (MDP) PUTERMAN1990, defined by the tuple with a finite vocabulary .
+The state space consists of all possible sequences of tokens up to the current time step, which includes the input sequence in the source language, as well as the target language tokens generated so far.
 Initially, the state corresponds to the input sentence in the source language, , where each token .
@@
 For , the state transition function defines the probability of transitioning from one state to another by appending a chosen token to the current translation, and the reward function assigns a real-valued reward to each transition , where and , based on the quality of the generated translation sequence.
-Explicitly, the reward function can be written as:
-Sequence-level rewards are provided only once per translation and evaluate the entire output at once.
-Token-level rewards, on the other hand, give feedback for each generated token.
-The discount factor is used to weigh future rewards, with typically chosen in MT to ensure that all rewards are valued equally, allowing the optimization of the entire sequence of tokens in the translation rather than focusing on just the initial tokens.
+Conceptually, the reward function is defined as a mapping from a hypothesis to a score, i.e., .
+In practice, many MT metrics additionally condition on the source and/or the reference, which we make explicit as .
+Formally, the reward function can be written as:
+Sentence-level rewards are provided only once per translation and evaluate the entire output at once.
+Token-level rewards, on the other hand, give feedback for each generated token.The discount factor is used to weigh future rewards, with typically chosen in MT to ensure that all rewards are valued equally, allowing the optimization of the entire sequence of tokens in the translation rather than focusing on just the initial tokens.
 Finally, the goal is to maximize the expected cumulative reward over trajectories sampled from the .
@@
 Policy Gradient Algorithms
-To optimize the above objective, we can use policy gradient methods: REINFORCE Williams (1992); Ranzato et al.
-(2015), a vanilla policy gradient method, optimizes translation by sampling hypotheses , scoring them with a reward obtained from an MT metric , and updating the model to maximize expected rewards, as shown in Equation 2.
+To optimize the above objective, we can use policy gradient methods: REINFORCE WILLIAMS1992; RANZATO2016, a vanilla policy gradient method, optimizes translation by sampling hypotheses , scoring them with a reward obtained from an MT metric , and updating the model to maximize expected rewards, as shown in Equation 2.
 Despite its simplicity, it often struggles with high variance and instability.
-Proximal Policy Optimization (PPO) Schulman et al.
-(2017) mitigates this by using a clipped surrogate objective (Equation 3) to keep policy updates stable within a margin , efficient and employs Generalized Advantage Estimation (GAE) Schulman et al.
-(2015) to compute advantages using rewards and value function .
-While PPO performs well across various tasks, simpler methods like REINFORCE can sometimes rival or surpass it Ahmadian et al.
-(2024).
+Proximal Policy Optimization (PPO) PPO mitigates this by using a clipped surrogate objective (Equation 3) to keep policy updates stable within a margin , efficient and employs Generalized Advantage Estimation (GAE) schulman2018highdimensionalcontinuouscontrolusing to compute advantages using rewards and value function .
+While PPO performs well across various tasks, simpler methods like REINFORCE can sometimes rival or surpass it ahmadian2024basics.
 Both are evaluated in our experiments (§5.2).
@@
 Advancements and Challenges in Sentence-level Feedback.
-Incorporating human feedback as rewards and optimizing language models with RL methods effectively aligns them with human preferences Ouyang et al.
-(2022), often surpassing MLE.
-A notable example in translation tasks is Minimum Risk Training (MRT) Shen et al.
-(2016), which minimizes expected risk based on evaluation metrics to directly improve translation quality.
-Recent advances in NMT build on this idea by refining training with feedback from metrics or human evaluations, incorporating alignment techniques and RL methods Nguyen et al.
-(2017); Kreutzer et al.
-(2018b); Wu et al.
-(2018); Kiegeland and Kreutzer (2021); Xu et al.
-(2023); Agrawal et al.
-(2024); Zhu et al.
-(2024); He et al.
-(2024); Ramos et al.
-(2024).
-Despite these advancements, sentence-level feedback methods face persistent challenges such as sparse rewards, instability, and difficulty handling long sequences Wu et al.
-(2018).
-These issues hinder performance, generalization, and robust learning, even with multi-objective optimizations Wu et al.
-(2023); Jang et al.
-(2023).
-To address the limitations of sentence-level feedback, recent research has explored finer-grained rewards at the token level for tasks such as language model alignment Xia et al.
-(2024); Yoon et al.
-(2024); Guo et al.
-(2023); Cao et al.
-(2024), controllable text generation Li et al.
-(2024), query generation Ouyang et al.
-(2024), among others, but remain relatively underexplored in MT.
+Incorporating human feedback as rewards and optimizing language models with RL methods effectively aligns them with human preferences ouyang2022training, often surpassing MLE.
+A notable example in translation tasks is Minimum Risk Training (MRT) shen2016minimumrisktrainingneural, which minimizes expected risk based on evaluation metrics to directly improve translation quality.
+Recent advances in NMT build on this idea by refining training with feedback from metrics or human evaluations, incorporating alignment techniques and RL methods nguyen-etal-2017-reinforcement; kreutzer-etal-2018-reliability; wu-etal-2018-study; kiegeland-kreutzer-2021-revisiting; almapref; agrawal-etal-2024-modeling; Zhu2024; he-etal-2024-improving; ramos-aligning.
+Despite these advancements, sentence-level feedback methods face persistent challenges such as sparse rewards, instability, and difficulty handling long sequences wu-etal-2018-study.
+These issues hinder performance, generalization, and robust learning, even with multi-objective optimizations wu2023finegrained; jang2023personalizedsoupspersonalizedlarge.
+To address the limitations of sentence-level feedback, recent research has explored finer-grained rewards at the token level for tasks such as language model alignment xia-etal-2024-inverse; yoon-etal-2024-tlcr; 2023arXiv231104072G; cao-etal-2024-enhancing, controllable text generation li-etal-2024-reinforcement, query generation 2024arXiv241100722O, among others, but remain relatively underexplored in MT.
 Token-level Feedback and Reward Modeling for MT.
-Previous approaches to token-level reward modeling often relied on binary error markings generated by humans Kreutzer et al.
-(2020); Domingo et al.
-(2017) or simulated it by comparing model predictions with reference translations based on heuristic methods Petrushkov et al.
-(2018).
+Previous approaches to token-level reward modeling often relied on binary error markings generated by humans kreutzer-etal-2020-correct; domingo2017segment or simulated it by comparing model predictions with reference translations based on heuristic methods petrushkov-etal-2018-learning.
 While effective, these methods provide limited feedback due to their binary nature and require costly human annotation, making them less practical for scalable solutions.
-Other approaches have employed reward shaping techniques Ng et al.
-(1999); Wu et al.
-(2018); Goyal et al.
-(2019); Devidze et al.
-(2022), incorporating intermediate rewards along with BLEU Papineni et al.
-(2002) as the reward function.
+Other approaches have employed reward shaping techniques ng1999policy; wu-etal-2018-study; RS_GOYAL; RS_RATI, incorporating intermediate rewards along with BLEU bleu as the reward function.
 However, partial BLEU or token-level BLEU are less effective for fine-grained reward modeling, as they depend on exact -gram matching and fail to capture meaningful semantic differences and context.
 Consequently, these methods, while valuable, are limited in their granularity and fail to address the severity of errors introduced at the token level.
-To address these limitations, in this work, we instead use token-level rewards obtained from state-of-the-art evaluation metrics that predict error spans and error severity levels to improve the learning and translation quality of MT systems.
 4 Approach
 In this section, we present our method for incorporating token-level rewards into RL training for machine translation (MT).
-We first detail how error assessments are mapped to token-level reward scores and then describe our adaptation of the REINFORCE and PPO learning objectives at the token level to enable learning from this fine-grained feedback.
+To address the limitations of prior approaches, such as binary feedback or coarse sentence-level scores, we use token-level rewards derived from state-of-the-art evaluation metrics that predict error spans and severity levels.
+These fine-grained signals are then used to guide learning through adaptations of REINFORCE and PPO objectives at the token level, enabling more effective and stable training of MT systems.
 Token-level Reward Modeling.
@@
 minor, major, critical) for a hypothesis given source and optionally a gold reference, we assign numerical weights to each token within an error span according to a severity mapping as defined below:
-We use the evaluation metric, xCOMET as as it was shown to achieve the best correlation with human judgments and was the winning submission for the WMT23 Metrics Shared Task Freitag et al.
-(2023).
-The severity weights from xCOMET adhere to the MQM framework Lommel et al.
-(2014), which classifies translation issues into categories such as fluency, adequacy, grammar, and style.
+We use the evaluation metric, xCOMET as as it was shown to achieve the best correlation with human judgments and was the winning submission for the WMT23 Metrics Shared Task freitag-etal-2023-results.
+The severity weights from xCOMET adhere to the MQM framework lommel-mqm, which classifies translation issues into categories such as fluency, adequacy, grammar, and style.
 Each token within an error span is assigned the same severity weight (see Figure 1).
@@
 Tokenization-Agnostic Reward Assignment.
-MT as well as evaluation models often rely on subword-level tokenization methods, such as Byte-Pair Encoding (Gage, 1994, BPE) or SentencePiece Kudo and Richardson (2018), where a single word can be split into multiple tokens.
-For a detokenized hypothesis generated by the MT system, identifies detokenized error spans corresponding to the errors.
-To generate token-level reward assignments for the tokenized hypothesis, , we assign the same reward to all subword tokens that belong to the same parent word.
-This is done by identifying the boundaries of the original word and ensuring that all tokens within those boundaries share the same reward.
-In this way, token-level rewards remain robust and independent of the tokenization differences between the generation and the evaluation models.
+MT systems and evaluation models typically use subword-level tokenization methods such as Byte-Pair Encoding (BPE, BPE) or SentencePiece kudo-richardson-2018-sentencepiece, where words can be split into multiple subword tokens and token boundaries may not align with natural word boundaries.
+Given a detokenized hypothesis from the MT system, our evaluation model produces error spans defined at the character level.
+To assign rewards at the token level for the tokenized hypothesis , we first re-tokenize the detokenized hypothesis using the same tokenizer applied during model training.
+This allows us to obtain precise character offsets for each subword token.
+We then align tokens to the character-level error spans by checking for overlap: any token whose character span overlaps with an error span inherits the corresponding error severity.
+This alignment avoids relying on explicit word boundaries or whitespace segmentation, making the reward assignment robust to different tokenization schemes – including those that generate cross-word subword units – and applicable across languages with or without explicit word boundaries.
+By grounding token-level rewards in character-level overlap rather than word-based grouping, our method ensures consistency and generalizability across tokenization models and languages.
+Finally, if a token overlaps multiple spans, it is assigned the worst severity – critical major minor correct – avoiding averaging or length-weighting to remain tokenizer-agnostic.
+Formally, for a token with overlaps , if is empty, .
+This severity is then mapped to its numeric reward (Table 4).
+The full algorithm is provided in Appendix A.
 Token-level Policy Refinement.
@@
 The policy ratio captures the change in policy for each token relative to the previous policy.
-The token-level advantage measures the improvement in performance over a baseline for each token and is used to scale updates based on expected reward improvements, as described in Schulman et al.
-(2015, 2017).
+The token-level advantage is estimated using Generalized Advantage Estimation (GAE) (schulman2018highdimensionalcontinuouscontrolusing), which balances bias and variance by mixing temporal-difference (TD) errors across multiple steps.
+The advantage at time step is computed as:
+where is the temporal-difference (TD) error, is the GAE parameter, is the reward at time step , and is a learned value function that estimates the expected return from the state defined by the input and the generated prefix .
+As explained earlier, we consider for our use case, and therefore omit the discount factor for notational simplicity.
 To ensure stable training, we apply clipping to limit the extent of policy updates, preventing large, unstable shifts.
@@
 Models.
-We use three state-of-the-art models: a standard encoder-decoder MT model, NLLB111facebook/nllb-200-1.3B NLLB Team et al.
-(2022), and two LLM-based MT systems, Tower222Unbabel/TowerInstruct-Mistral-7B Alves et al.
-(2024) and Gemma333https://huggingface.co/google/gemma-2-9b-it Gemma Team et al.
-(2024).
+We use three state-of-the-art models: a standard encoder-decoder MT model, NLLB flores, and two LLM-based MT systems, Tower tower and Gemma gemmateam2024gemma2improvingopen.
 While NLLB and Tower are dedicated MT models optimized for translation tasks, Gemma is an LLM that exhibits strong multilingual capabilities.
@@
 Data.
-We use the following training datasets in our experiments: (1) The IWSLT2017 dataset Cettolo et al.
-(2017), with k examples for English-French (ENFR), supports rapid experimentation and frequent training iterations.
-(2) The WMT18 dataset Bojar et al.
-(2018) contains M examples for English-German (ENDE).
+We use the following training datasets in our experiments: (1) The IWSLT2017 dataset iwslt2017, with k examples for English-French (ENFR), supports rapid experimentation and frequent training iterations.
+(2) The WMT18 dataset WMT18 contains M examples for English-German (ENDE).
 We train NLLB with both datasets and the LLM-based models with (2).
 Training stops once rewards stabilize, so not all examples are used.
-We evaluate NLLB models using the IWSLT17 ENFR and WMT18 ENDE test sets.
-To standardize comparison across MT systems (NLLB and LLM-based MT systems), we also evaluate all models on the WMT24 dataset Kocmi et al.
-(2024), addressing concerns about data contamination, as Tower training included the WMT18 test set.
+We evaluate NLLB models using their respective test splits: IWSLT17 (ENFR) and WMT18 (ENDE).
+To standardize comparison across MT systems (NLLB and LLM-based MT systems), we also evaluate all models on the WMT24 dataset WMT24, addressing concerns about data contamination, as Tower training included the WMT18 test set.
 Evaluation.
 We assess translation quality using a comprehensive suite of well-established evaluation metrics.
-These include lexical reference-based metrics, such as BLEU Papineni et al.
-(2002) and ChrF Popović (2015); neural reference-based metrics, including COMET22 Rei et al.
-(2022), xCOMET Guerreiro et al.
-(2024), and BLEURT Sellam et al.
-(2020); and a neural reference-free metric, CometKiwi-23 Rei et al.
-(2023).
+These include lexical reference-based metrics, such as BLEU bleu and ChrF chrf; neural reference-based metrics, including COMET22 comet22, xCOMET xcomet, and BLEURT bleurt; and a neural reference-free metric, CometKiwi-23 cometkiwi.
 Lexical metrics focus on word overlap and -gram matching, while neural metrics evaluate translations in terms of semantic coherence and contextual quality.
@@
 We apply significance testing at a confidence threshold of .
-For segment-level metrics, such as COMET-22, we test at the segment level, but for corpus-level metrics, such as BLEU and ChrF, we apply bootstrapping with samples of size Koehn (2004).
-Performance clusters are formed based on statistically significant gaps, and final rankings are derived by averaging the cluster scores across all languages Colombo et al.
-(2022); Freitag et al.
-(2023).
-In addition to automated metrics, we conduct human evaluations to gain deeper insights into translation performance.
-Furthermore, significance testing is performed to validate improvements observed across different models and approaches, ensuring the reliability of our findings.
+For segment-level metrics, such as COMET-22, we test at the segment level, but for corpus-level metrics, such as BLEU and ChrF, we apply bootstrapping with samples of size koehn-2004-statistical.
+Performance clusters are formed based on statistically significant gaps, and final rankings are derived by averaging the cluster scores across all languages colombo2022bestsystemsnewperspectives; freitag-etal-2023-results.
+In addition to automated metrics, we conduct human evaluations with two professional annotators, reporting inter-annotator agreement (Pearson’s and Spearman’s ) and confidence intervals across length buckets to assess the reliability of model differences.
 Reward Models.
-We utilize two reward models based on xCOMET444https://huggingface.co/Unbabel/XCOMET-XL.
+We utilize two reward models based on xCOMET.
 The MQM-derived reward signal, referred to as xCOMET-MQM, is generated from error span predictions that identify and classify translation errors by severity.
@@
 Training Configurations.
-We finetune NLLB, Tower and Gemma models using MLE and RL methods detailed in §3 with the following configurations:
+We finetune NLLB, Tower and Gemma models using MLE and RL methods detailed in Section 3 with the following configurations:
 SFT: a baseline model supervised finetuned on the parallel data using MLE (1).
 sRL: We compare using sentence-level xCOMET with BLEU. The learning algorithm used is PPO (3), a current state-of-the-art alignment method for MT.
-CPO Xu et al.
-(2024): a state-of-the-art preference optimization learning method for MT, offering a more efficient variant of DPO Rafailov et al.
-(2023).
-We construct the preference dataset by generating multiple outputs from the MT model using the training datasets,555We generate 16 samples with the value of top_p top_k set to 0.9 and 50 respectively.
+CPO CPO: a state-of-the-art preference optimization learning method for MT, offering a more efficient variant of DPO DPO.
+We construct the preference dataset by generating multiple outputs from the MT model using the training datasets,111We generate 16 samples with the value of top_p top_k set to 0.9 and 50 respectively.
 and then induce preferences using the xCOMET metric, comparing these outputs to human-written references.
@@
 We use the same hyperparameter settings for NLLB, Tower, and Gemma.
-We use HuggingFace’s Transformers library Wolf et al.
-(2020) and the Transformers Reinforcement Learning (TRL) library 666https://github.com/huggingface/trl to facilitate RL training.
-We perform MLE training with Adam Kingma and Ba (2014) as the optimization algorithm, learning rate decay starting from and early stopping.
+We use HuggingFace’s Transformers library wolf-etal-2020-transformers and the Transformers Reinforcement Learning (TRL) library to facilitate RL training.
+We perform MLE training with Adam kingma2017adam as the optimization algorithm, learning rate decay starting from and early stopping.
 We use PPO with a learning rate of , set as , trajectory limit set as .
 Mini-batch updates are performed with a batch size of over PPO epochs.
-The translation prompt for LLM-based MT systems is shown below:
+The translation prompt for LLM-based MT systems is shown in Table 1.
 | Translate the following text from {source_lang} into {target_lang}. |
@@
 tRL consistently outperforms SFT and sRL methods across neural metrics.
-For all language directions reported in Table 2, “tRL w/ xCOMET-MQM“ outperforms SFT and its sentence-level counterpart “sRL w/ xCOMET-MQM“ across all neural metrics considered.
+For all translation directions reported in Table 2, “tRL w/ xCOMET-MQM“ outperforms SFT and its sentence-level counterpart “sRL w/ xCOMET-MQM“ across all neural metrics considered.
 SFT significantly improves translation quality by tailoring the pre-trained MT model to the specific target language pairs.
@@
 When comparing the sRL and tRL methods, we observe that sRL leads to moderate improvements over SFT, while the gains obtained by tRL are more substantial, particularly when assessed with advanced neural metrics.
-Although the ChrF scores for tRL models trained with xCOMET-MQM are lower, possibly due to the focus on optimizing token-level features over character-level accuracy, all other metrics indicate significant improvements, suggesting the overall translation quality is better.
+Although the chrF scores for tRL models trained with xCOMET-MQM are lower on IWSLT2017, this is likely due to a mismatch between the reward signal and the evaluation metric: token-level rewards optimize semantic quality, not character-level overlap.
+This can lead to fluent, accurate translations that diverge lexically from references, reducing chrF despite improved overall quality.
+Neural metrics, which are more robust to surface-level variation and better aligned with human judgments freitag-etal-2022-results; freitag-etal-2023-results, as well as human evaluation (see Section 5.4) – the gold standard for assessing translation quality – consistently show improvements with our tRL approach.
+Appendix C presents a focused quantitative and qualitative analysis of cases in which chrF decreases while xCOMET improves.
 | Model | Metrics | |||||
@@
 tRL improves translation quality for LLM-based MT systems, Tower and Gemma.
-Our severity-based fine-grained reward optimization mechanism significantly improves translation quality across all automatic evaluation metrics as shown in Table 3.
-These findings highlight that tRL not only improves the quality of state-of-the-art MT models but can also significantly boost stronger MT-LLMs, demonstrating its broad applicability and potential for advancing multilingual translation systems.
+Our severity-based, fine-grained mechanism significantly improves translation quality across all automatic evaluation metrics, as shown in Table 3.
+These findings highlight that tRL not only improves the quality of state-of-the-art MT models but can also significantly boost stronger LLM-based MT systems, demonstrating its broad applicability and potential for advancing multilingual MT systems.
 On-policy PPO results in better translation quality than RL-free method, CPO.
@@
 Given that xCOMET exhibits a strong correlation with human judgments, it proves to be an essential tool for guiding MT models toward higher translation quality.
-In contrast, traditional metrics like BLEU, based on -gram overlap, can fall short in aligning with human judgments as they do not capture contextual nuances and semantic understanding Freitag et al.
-(2022).
+In contrast, traditional metrics like BLEU, based on -gram overlap, can fall short in aligning with human judgments as they do not capture contextual nuances and semantic understanding freitag-etal-2022-results.
 Consequently, BLEU performs less effectively compared to neural metrics like xCOMET in this setup which use contextual embeddings.
@@
 Smooth transition severity map to result in better translation quality.
-In contrast, abrupt changes in the reward signal can destabilize learning, leading to inconsistent training, oscillations, or convergence to suboptimal policies Ranzato et al.
-(2015); Sutton and Barto (2018).
+In contrast, abrupt changes in the reward signal can destabilize learning, leading to inconsistent training, oscillations, or convergence to suboptimal policies RANZATO2016; sutton2018reinforcement.
 Additionally, we find that the binary severity map, which ignores the severity of errors, provides less informative feedback to the model, resulting in slightly lower performance than maps that offer more nuanced assessments.
@@
 tRL improves training stability over sRL.
-Figure 3 shows the evolution of mean rewards during training for sRL and tRL methods across the two datasets for the NLLB system.
-As can be observed, tRL training shows a more stable and consistently increasing reward evolution, which is crucial for the training process as it ensures consistent improvements and reduces the risk of overfitting or fluctuations that could degrade the MT model’s performance.
-tRL improves translation quality for longer sequences.
-Building on our hypothesis that tRL is particularly effective for longer sequences, we show the COMET22 scores for the WMT24 EN DE dataset, grouped by sequence length, across different training methods, including NLLB and Tower, in Figure 4.
-tRL consistently outperforms other methods, especially for longer sequences, which we attribute to its ability to capture localized and fine-grained signals.
-These results further support our findings, showing that tRL remains the most robust method, with scores dropping less than other approaches as the source sentence length increases.
-This reinforces the effectiveness of tRL in handling complex, longer sentence structures.
+Figure 3 shows the evolution of mean rewards during training for sRL and tRL across the two datasets for the NLLB system.
+As observed, tRL training exhibits a more stable and consistently increasing reward trajectory, which is crucial for ensuring steady improvements and reducing the risk of performance-degrading fluctuations or overfitting.
 | Model | Metrics | |||||
@@
 | + tPPO w/ xCOMET-MQM | 46.58 | 60.07 | 87.17 | 92.17 | 75.58 | 72.80 |
+tRL improves translation quality for longer sequences.
+Building on our hypothesis that tRL is particularly effective for longer sequences, we present COMET22 scores for the WMT24 ENDE dataset in Figure 4, grouped by source sequence length and comparing different training methods, including NLLB and Tower.
+The figure also shows the distribution of source sequence lengths in the training and test data.
+Notably, the training data is skewed toward shorter inputs—a common characteristic of large MT corpora—whereas the WMT24 test set includes a broader distribution with a higher proportion of long sequences.
+This discrepancy highlights the need for models that generalize well to longer inputs.
+In this context, "longer sequences" refers to short paragraphs rather than single sentences.
+These are not document-level inputs and remain within the 512-token limit of xCOMET, ensuring that the reward model processes them without truncation during training.
+tRL consistently outperforms other training methods, especially on longer sequences.
+We attribute this to its ability to capture localized, fine-grained reward signals during training.
+These findings further support our earlier results: tRL demonstrates the most robust performance, with smaller performance drops as the source sentence length increases, confirming its strength in handling complex sentence structures.
+For completeness, we also evaluate in Appendix B a hybrid model that applies sRL to short inputs and tRL to long ones.
+This approach outperforms sRL but remains inferior to tRL.
 REINFORCE and PPO are suitable methods for training MT systems.
-We compare REINFORCE and PPO for RL-based MT with xCOMET as reward model, evaluating their impact on translation quality (Table 6).
+We compare REINFORCE and PPO for RL-based MT with xCOMET as the reward model, evaluating their impact on translation quality (Table 6).
 Both methods are effective, but PPO achieves superior overall metric scores due to features such as objective clipping and KL divergence control, which enhance training stability.
 However, REINFORCE remains a strong alternative for simpler implementations that aim to achieve competitive performance.
-| Reward Model | BLEU | xCOMET |
-| Latency | ||
-| Throughput | ||
-| Quality |
 Efficiency and Quality Tradeoffs in Token-Level Reward Computation.
-tRL using xCOMET offers higher quality, as measured by COMET22, than tRL with BLEU but comes with increased computational costs, including higher latency and lower throughput due to its larger pre-trained encoder, as reported in Table 7.
-Despite the slower processing speeds, xCOMET’s superior reward quality and resulting higher model quality are particularly valuable in token-level feedback scenarios, where high-quality rewards are essential, and the additional computational cost is manageable with GPU acceleration.
-Furthermore, advances in compute efficiency for transformer-based architectures, such as quantization Jacob et al.
-(2017), FlashAttention Dao et al.
-(2022); Dao (2023), and distillation Hinton et al.
-(2015), can help reduce the computational load, making xCOMET more practical for broader applications.
+Using xCOMET as a reward model for tRL yields higher-quality translations than BLEU, but it also incurs increased computational costs, as detailed in Table 7.
+Due to its larger pre-trained encoder, xCOMET exhibits significantly higher latency (average seconds per token) and lower throughput (tokens per second).
+It is worth noting that throughput values can appear higher than latency alone might suggest, as this metric benefits from amortized per-call overheads and batching.
+The quality improvement, measured by COMET22, reflects the performance of our best model (Tower) on WMT24 ENDE after fine-grained optimization with each reward model.
+Despite the slower processing, xCOMET’s superior reward quality is particularly valuable in token-level feedback scenarios where high-quality rewards are essential and the additional computational cost is manageable with GPU acceleration.
+Furthermore, advances in computational efficiency for transformer architectures—such as quantization quant, FlashAttention fa; fa2, and distillation dist—can help mitigate this computational load, making xCOMET more practical for broader applications.
+| Latency | Throughput | Quality | |
+| BLEU | |||
+| xCOMET |
 5.4 Human Evaluation
 Setup.
-For our human evaluation, we used Direct Assessments (Graham et al., 2013, DAs) to score translations on a scale from 0 to 100, following the standard WMT human evaluation methodology.
+For our human evaluation, we used Direct Assessments (graham-etal-2013-continuous, DAs) to score translations on a scale from 0 to 100, following the standard WMT human evaluation methodology.
 We evaluate 200 randomly chosen instances from the WMT18 EN DE dataset.
-Two professional translators, both native speakers of the target language, assesse the references and NLLB translations using the following methods: SFT, CPO with xCOMET, tRL with xCOMET and our proposed severity mapping, and sRL with xCOMET.
-While WMT18 EN DE allows for easy comparison between several methods due to shorter sequences, we conducte a second human evaluation on the WMT24 EN DE dataset to validate our empirical finding that tRL benefits longer input sequences.
+Two professional translators, both native speakers of the target language, assess the references and NLLB translations using the following methods: SFT, CPO with xCOMET, tRL with xCOMET, our proposed severity mapping, and sRL with xCOMET.
+While WMT18 EN DE allows for easy comparison between several methods due to shorter sequences, we conducted a second human evaluation on the WMT24 EN DE dataset to validate our empirical finding that tRL benefits longer input sequences.
 For this setting, we directly compare sRL with xCOMET and tRL with xCOMET in a pairwise setting using outputs from the Tower model.
 To ensure adequate coverage across different sequence lengths, we performed stratified sampling based on source length, using bins [0, 100, 250, 500, 1000] with 50 instances per bin.
-| Model | Ann1 | Ann2 |
-| CPO | 59.5 | 53.2 |
-| SFT | 64.5 | 54.2 |
-| tRL | 66.9 | 56.1 |
-| sRL | 66.8 | 57.6 |
-| Reference | 76.0 | 60.0 |
+| CPO | SFT | tRL | sRL | Reference | |
+| Ann1 | 59.5 | 64.5 | 66.9 | 66.8 | 76.0 |
+| Ann2 | 53.2 | 54.2 | 56.1 | 57.6 | 60.0 |
 Findings.
-As shown in Table 8, both sRL and tRL models consistently outperform SFT and CPO, demonstrating their advantage in enhancing translation quality.
-On average, human evaluation scores for tRL and sRL are similar.
-On the WMT24 dataset, tRL achieves an average DA score of 82.6, 2.3 points higher than sRL.
-We additionally show the distribution of DA scores bucketed by length in Figure 5.
-Across all buckets, tRL outperforms sRL, with more significant (based on error bars overlap) gains for longer sentences.
-These findings are also consistent with our automatic evaluation, including the ablation analysis (§ 5.3) and clearly demonstrate that tRL enhances stability and translation quality, particularly for longer sentences.
+As shown in Table 8, both sRL and tRL models consistently outperform SFT and CPO, demonstrating their advantage in translation quality.
+On the WMT24 dataset, tRL achieves an average DA score of , points higher than sRL, with consistent gains across sentence-length buckets, as shown in Figure 5; longer sentences exhibit more significant improvements based on error bar overlap.
+Human evaluations were conducted by two professional annotators, and inter-annotator agreement is moderate-to-high (Pearson’s , Spearman’s ), indicating reliable scoring.
+These results align with our automatic evaluation, including the ablation analysis, which collectively shows that tRL enhances stability and translation quality, particularly for longer sentences.
 6 Conclusion
@@
 This work was supported by the Portuguese Recovery and Resilience Plan through project C645008882-00000055 (Center for Responsible AI), by the EU’s Horizon Europe Research and Innovation Actions (UTTER, contract 101070631), by the project DECOLLAGE (ERC-2022-CoG 101088763), and by Fundação para a Ciência e Tecnologia through contract UIDB/50008/2020.
+Appendix A Details of the Severity Assignment Algorithm
+In this section, we provide a detailed description of the severity assignment algorithm used in our approach, focusing on the case of tokens overlapping annotated error spans.
+Our implementation resolves multiple overlapping error spans by assigning the token the worst severity among all overlapping spans.
+This choice aligns with MQM annotation practices, where the most severe issue in a region governs its quality classification.
+Formally, if a token overlaps spans labeled minor, major, and critical, the token is assigned critical.
+We avoid averaging or length-weighted schemes to remain fully tokenizer-agnostic.
+A token is considered affected by a span if it overlaps with it in any way, not only if it is fully contained.
+This avoids mismatches in cases where tokens are longer than the annotated spans.
+Formally, let
+be the set of tokens,
+be the set of annotated error spans,
+denote the severity of span .
+The severity assigned to token is defined as:
+Example.
+Suppose the text is tokenized as a single token “abc”, and error spans are defined as (minor), (major), and (critical).
+Since the token “abc” overlaps with all three spans, its assigned severity is critical.
+Appendix B Details of the Hybrid sRL–tRL Experiment
+| Method | Metrics | |||||
+| BLEU | ChrF | COMET22 | xCOMET | BLEURT | CometKiwi-23 | |
+| Tower | 42.77 | 62.60 | 71.80 | 88.50 | 68.59 | 67.20 |
+| + SFT | 45.14 | 63.30 | 72.18 | 88.90 | 69.18 | 67.30 |
+| + sRL w/ xCOMET-MQM | 45.58 | 63.76 | 72.41 | 90.20 | 70.55 | 69.70 |
+| + tRL w/ xCOMET-MQM | 46.92 | 65.63 | 74.66 | 91.90 | 71.80 | 71.20 |
+| + hRL w/ xCOMET-MQM | 45.96 | 65.07 | 73.20 | 90.57 | 71.30 | 69.34 |
+To examine whether the improvements of tRL are primarily driven by long-sequence behavior, we evaluate a hybrid reinforcement learning (hRL) approach that applies sentence-level RL (sRL) to short inputs and token-level RL (tRL) to long ones.
+Short sentences are defined as those below the average source length in the training data, and long sentences as those above it.
+This experiment follows our best-performing configuration: the Tower model with xCOMET-MQM used as the reward signal.
+Table 9 reports the results alongside the relevant baselines from Table 3.
+The hRL model improves over sRL on most metrics but consistently falls short of tRL, indicating that tRL provides the strongest training signal across sentence lengths without the added complexity of a hybrid setup.
+Appendix C Analysis of chrF Drop Cases
+This section provides a focused quantitative and qualitative analysis of translations where chrF decreases but xCOMET improves under token-level RL (tRL).
+The goal is to examine whether the observed chrF drop corresponds to genuine translation degradation or reflects a metric mismatch.
+We leverage available human evaluation data for WMT24 ENDE, using Direct Assessment (DA) scores as reliable indicators of translation quality.
+This analysis stems from the observation that, in some settings, tRL outputs show a noticeable drop in chrF (up to 7 points on IWSLT2017) while simultaneously improving xCOMET and other quality metrics.
+This raised concerns that such a drop might reflect lexical imprecision or other undesirable artifacts.
+To investigate, we conducted a detailed error analysis focusing on cases with the largest discrepancies between chrF and xCOMET.
+Examining these high-divergence examples provides the clearest insight into whether chrF drops correspond to real quality issues or simply reflect a mismatch between metrics.
+C.1 Quantitative Analysis
+Figure 6 summarizes the extreme discrepancy cases for ENDE translations.
+The plot shows the average DA score for the top-N cases with the largest chrF–xCOMET discrepancies.
+Even with the observed chrF drops and corresponding xCOMET increases, tRL consistently achieves higher human DA scores than sRL, indicating that translation quality is not compromised.
+This strongly suggests that the apparent chrF decline is a metric artifact rather than a genuine degradation in translation quality.
+C.2 Qualitative Analysis
+Table 10 presents representative examples of translations with chrF drops but higher xCOMET and DA scores.
+These cases illustrate how lexical divergence from the reference can lower chrF while yielding translations that are more fluent, semantically accurate, and preferred by human evaluators.
+| Example 1 | |
+| Source | I like both the glasses you posted, these are a really nice colour…makes me want new ones |
+| Reference | Ich mag beide Brillen, die du gepostet hast, die haben eine wirklich nette Farbe … weckt bei mir den Wunsch, neue zu kaufen |
+| sRL translation | Ich mag beide Brillen, die du gepostet hast, diese sind wirklich eine schöne Farbe… Ich würde gerne neue haben |
+| tRL translation | Mir gefallen beide Brillen, die du gepostet hast, sie haben eine wirklich schöne Farbe… jetzt will ich auch neue |
+| Scores | ChrF 63.4 59.7 xCOMET: 96.8 99.3 DA: 75.0 84.0 |
+| Example 2 | |
+| Source | I don’t have any telemetry for the battery. This is something I might add in a future revision. |
+| Reference | Ich habe keine Telemetrie für den Akku. So etwas könnte ich in Zukunft noch einbauen. |
+| sRL translation | Ich habe keine Telemetrie für den Akku. Das ist etwas, was ich in einer zukünftigen Überarbeitung hinzufügen könnte. |
+| tRL translation | Ich habe keine Telemetrie-Daten für den Akku. Das ist etwas, was ich vielleicht in einer zukünftigen Version hinzufügen könnte. |
+| Scores | ChrF 58.4 53.1 xCOMET: 97.2 98.2 DA: 84.0 94.0 |
+| Example 3 | |
+| Source | “Which notebook is that?” Ivory asked, sitting down next to Kari on her bed. |
+| Reference | „Welches Notizbuch ist das?“, fragte Ivory und setzte sich neben Kari auf deren Bett. |
+| sRL translation | „Welches Notizbuch ist das?“ fragte Ivory und setzte sich neben Kari aufs Bett. |
+| tRL translation | „Welches Notizbuch ist das?“, fragte Ivory, während sie sich neben Kari auf das Bett setzte. |
+| Scores | ChrF 85.3 80.4 xCOMET: 99.3 99.5 DA: 84.0 93.0 |
+| Example 4 | |
+| Source | I’m splurging on a new set of frames, these red ones I reeeeally like. |
+| Reference | Ich verschwende mein Geld für eine neue Brillenfassung, diese rote mag ich seeehr. |
+| sRL translation | Ich habe mir eine neue Brille gekauft, diese rote gefällt mir richtig gut. |
+| tRL translation | Ich gönne mir ein neues Brillengestell, diese roten gefallen mir ganz besonders. |
+| Scores | ChrF 37.8 33.8 xCOMET: 95.2 95.5 DA: 76.0 92.0 |
+| Example 5 | |
+| Source | “Well my buddys from boot camp here did” Cohren motioned to Harris, Craith and Ravik. “But then again, their squad and platoon leader’s, they know to wake up before their units. |
+| Reference | „Nun, bei meinem Kumpels aus dem Bootcamp ist das bereits der Fall“, Cohren winkte Harris, Craith und Ravik zu.
+„Aber auch ihre Truppen- und Zugführer wissen, dass sie vor ihren Einheiten aufwachen müssen.“ |
+| sRL translation | „Nun ja, meine Kameraden vom Bootcamp hier haben es getan“, sagte Cohren zu Harris, Craith und Ravik.
+„Aber ihr Zug- und Platoonführer wissen, dass sie vor ihren Einheiten aufwachen müssen.
+|
+| tRL translation | „Nun, meine Kameraden vom Bootcamp haben es getan“, sagte Cohren und deutete dabei auf Harris, Craith und Ravik.
+„Aber sie und ihre Zugführer wissen ja auch, dass sie vor ihren Einheiten aufstehen müssen.“ |
+| Scores | ChrF 57.8 56.2 xCOMET: 81.9 85.1 DA: 74.0 85.0 |